• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eric

    Log entry time 00:36:03 on October21,2002

    Entry number 88574

    keyword=Shift summary


    New crew on shift: Martin Epstein, Paul Ulmer, Eric Voutier
    ***************

    17:05 Call MCC to have an update: the BCM team change some device bu they
    dont know if this will solve the problem; they will send shortly the beam into
    the hall to check the BCM.

    17:20 MCC want to send low CW current in the line for checks; move to optics
    target for safety... 5mn later they call back for correction, they want to send the
    highest possible current (100 uA)

    18:25 Move to empty target since MCC wants to go to the highest possible
    current.

    18:50 MCC ready to send 20uA CW; moving to 15cm LH2 target.

    19:30 Start taking data on Ela_H02

    20:22 Start moving to Ela_H03

    21:13 Start data taking on Ela_H03

    23:00 While changing to the next kinematics, we experience problems with the
    left arm control screen: the momentum dont want to change... we called Javier.

    23:30 MCC calls the software on call to try to solve the problem... Javier and
    the software on call finally solve the problem: the software momentum limit
    was set to 3.16 GeV/c (right arm limit) which was forbidding us to go higher...
    problem is solved and we go gradually back to production.

    Run History
    -------------

    --# Ela_H02

    Run # 2593 .......... Good Run / 15cm LH2
    Run # 2594 .......... Good Run / 15cm LH2

    --# Ela_H03

    Run # 2595 .......... Good Run / 15cm LH2
    Run # 2596 .......... Good Run / 15cm LH2
    Run # 2597 .......... Good Run / Optics